# From Replication to Redesign: Exploring Pairwise Comparisons for LLM-Based Peer Review

### v1

@@ 1 Introduction
While LLMs offer new possibilities for peer review, simply replicating existing workflows may overlook a deeper opportunity. Historical analysis suggests that the conventional structure of peer review was not carefully engineered for optimal evaluation, but rather emerged as a pragmatic response to resource constraints. Studying the development of peer review in sociology, Merriman [29] observes that "the evolution of peer review is best understood as the product of continuous efforts to steward editors’ scarce attention while preserving an open submission policy that favors authors’ interests." This perspective highlights that many features of today’s review systems reflect historical compromises rather than principled design choices. As scholarly publishing expands globally and across disciplines [10], these inherited structures have come under increasing scrutiny. Issues like bias [19, 39, 41] and noisy ratings [25] have spurred debates on how to improve or reform peer review [14, 40, 43, 47].
Given their exceptional scalability, LLMs present an opportunity not merely to automate existing practices, but to fundamentally rethink the design of manuscript evaluation systems. Rather than replicating workflows optimized for human reviewers, it is crucial to reconsider what constitutes effective evaluation and to design AI systems that address the longstanding limitations of traditional peer review.
@@
In this work, we propose a shift from replication to redesign. We introduced and explored a novel mechanism that engages LLM agents in pairwise comparisons for more effective manuscript assessment. Rather than assigning absolute scores to individual papers, our approach directly contrasts pairs of submissions to determine which one better meets the multidimensional evaluation criteria. Each LLM agent is tasked with evaluating two manuscripts at a time, and by integrating the results from numerous pairwise comparisons, we can construct a robust overall assessment that more accurately reflects relative quality.
Our results indicate that with proper scaling, this pairwise mechanism demonstrates potential to identify papers of higher academic impact, significantly outperforming paper-level rating approach. However, our analysis also reveals important limitations: Papers preferred by our LLM-based pairwise ranking system tend to exhibit lower topic novelty and greater concentration among a limited number of high-prestige institutions. While these biases highlight critical challenges for the deployment of LLM-based review systems, we view this work as a first step toward exploring fundamentally new paradigms for scalable, equitable, and robust research evaluation.
@@ 2 Related Work
Peer review, while essential to scientific progress, faces numerous challenges in today’s academic landscape. Reviewers often struggle with overwhelming workloads, further compounded by the absence of meaningful incentives and the inequitable distribution of review requests [7, 28]. In addition, implicit biases can influence evaluation [41, 42, 45], with studies showing that factors like author affiliation or gender may affect acceptance rates [52]. As research output continues to accelerate globally [5], these systemic challenges demand innovative solutions to preserve the integrity and effectiveness of scientific evaluation.
@@
Recent advancements in Large Language Models (LLMs) have sparked considerable interest in their potential to transform academic peer review processes [20, 55]. While existing studies [9, 21, 34] have demonstrated that LLMs can provide valuable feedback for research papers, exhibiting substantial overlap with reviews written by human reviewers, they often struggle with assigning reasonable ratings and making sound decisions, which raises concerns about their reliability for evaluative tasks [23, 51]. For instance, Niu et al. [31] showed that ChatGPT tends to grant acceptances: out of 1558 full paper evaluations, ChatGPT suggested 1243 acceptances compared to only 315 rejections, resulting in an acceptance rate of approximately 79.8%. Similarly, Zhou et al. [53] highlighted that LLMs could generate reasonable aspect scores (e.g. recommendation, soundness, originality) from human reviews but failed when given only the research paper. Lu et al. [24] improved the base LLM’s decision-making process through prompting techniques such as self-reflection [38], few-shot examples [49] and response ensembling [48]. CycleReviewer [50] and DeepReview [54] enhanced LLM-based paper review via fine-tuning on high-quality review dataset. In contrast, Less effort is made to explore new paradigms that fundamentally rethink how LLMs can participate in the academic review process.
@@ 3.2 From Pairwise Comparisons to Rankings
    Pairwise Comparison: Let $\mathcal{S}=\{(i,j):i\neq j\text{ and }i,j\in[N]\}$ be all possible pairs of papers to compare. For each of $M$ agents, we randomly assign one pair $(i,j)\in\mathcal{S}$ to evaluate. Each agent then analyzes both papers and produces a binary preference judgment $y_{ij}\in\{0,1\}$, where $y_{ij}=1$ indicates that the agent prefers paper $i$ over paper $j$, and $y_{ij}=0$ indicates the opposite preference. The full prompt used for pairwise judgment is provided in Supp Figure 7. Denote the collection of all chosen pairs as $\mathcal{A}$, where $|\mathcal{A}|=M$. We collect all comparison results into an observation set $\mathcal{O}=\{(i,j,y_{ij}):(i,j)\in\mathcal{A}\}$, which contains triplets of the compared papers and their corresponding preference judgments from all agents. Since the size of $|\mathcal{S}|$ grows quadratically with $n$, we can only afford to assess a small fraction of all possible pairs, and hence it is challenging to recover the ranking under this scenario. Later in Section 4.2, we empirically examined how ranking quality scales with the number of LLM agents. The results reveal a clear scaling law, and we are able to recover a high-quality ranking with less than 2% samples of all possible pairs.
@@ 4.2 Agent Scaling Boosts the Performance of GPT Ranking System
Figure 2: Scaling of average citation counts for accepted papers by GPT ranking system with increasing pairwise comparisons. We introduce two variants of our GPT ranking system: a basic version without area control, and one with area control, which ensures that the distribution of accepted papers across primary research areas matches that of human peer review. These two variants are compared against three baselines: (1) human acceptance decisions, (2) random acceptance based on the conference acceptance rate, and (3) GPT rating system (acceptance depends on the rating synthesized by a meta-reviewer from three independent GPT reviewers). As the number of comparisons increases, both variants steadily improve and approach human-level performance. At maximum scale, the GPT ranking system without area control achieves 20.00 average citations, while the version with area control reaches 18.97 average citations. These compare favorably to human acceptance decisions (19.36 average citations), and significantly outperform both random acceptance (8.56 average citations) and the GPT rating system baseline (11.41 average citations).
@@ 4.3 Discriminative Capability of GPT ranking system in Research Evaluation
Figure 3: Comparison of mean citation counts across multiple AI conferences under two-fold decisions (e.g., accepted vs. rejected, main vs. findings track, oral vs. poster). The results consistently show that higher-tier papers (both GPT-selected and human-selected) receive substantially higher citation counts than lower-tier papers. This implies that our GPT ranking system can effectively distinguish more influential papers, performing comparably to human peer review system in identifying works likely to generate greater impact in the academic community. Error bars represent 95% confidence intervals. *P $<$ 0.05, **P $<$ 0.01, ***P $<$ 0.001, and ****P $<$ 0.0001.
@@ 4.6 Potential Bias against Novel Research Topics
Figure 5: Nearest neighbor distances of papers selected for top-tier acceptance by GPT ranking system versus human reviewers. Higher distances indicate greater novelty. Human-selected papers (blue) consistently show higher nearest neighbor distances than GPT-selected papers (orange) across all conferences.
@@ A.1 Limitations
In this study, we rely on citation counts as a proxy for academic impact. Although citation metrics are widely used and provide a quantifiable measure of influence, they are affected by numerous factors beyond quality, such as the visibility of research communities, prevailing research trends, and established reputations [22, 4, 46, 11]. More critically, papers selected by human reviewers inherently receive more visibility through conference presentations and proceedings. The fact that our GPT ranking system identified papers with comparable citation impact despite this disadvantage demonstrates a certain level of effectiveness in our approach. Nevertheless, this circularity problem makes it difficult to establish a truly independent measure of quality. Furthermore, while we exclusively used GPT-4o mini for our large-scale experiments due to its balance of performance and computational cost, we acknowledge that there are other diverse LLMs capable of academic assessment, and they might yield substantially different results. A more comprehensive evaluation across multiple models would provide better insight into the generalizability of our method and findings. Finally, while our current work constructs a single global ranking, future extensions could explore personalized or multi-objective evaluation systems that explicitly account for epistemic diversity and evolving community goals.
@@ C.1 Discriminative Capability of GPT ranking system in Research Evaluation
Figure 8: Comparison of mean citation counts across multiple AI conferences under fine-grained decisions in accepted papers. The results consistently show that higher-tier papers (both GPT-selected and human-selected) receive overall higher influence than lower-tier papers. Error bars represent 95% confidence intervals. *P $<$ 0.05, **P $<$ 0.01, ***P $<$ 0.001, and ****P $<$ 0.0001.

### v2

@@ 1 Introduction
While LLMs offer new possibilities for peer review, simply replicating existing workflows may overlook a deeper opportunity. Historical analysis suggests that the conventional structure of peer review was not carefully engineered for optimal evaluation, but rather emerged as a pragmatic response to resource constraints. Studying the development of peer review in sociology, Merriman [32] observes that "the evolution of peer review is best understood as the product of continuous efforts to steward editors’ scarce attention while preserving an open submission policy that favors authors’ interests."
This perspective highlights that many features of today’s review systems reflect historical compromises rather than principled design choices. As scholarly publishing expands globally and across disciplines [10], these inherited structures have come under increasing scrutiny. Issues like bias [20, 44, 46] and noisy ratings [28] have spurred debates on how to improve or reform peer review [14, 45, 48, 52].
Given their exceptional scalability, LLMs present an opportunity not merely to automate existing practices, but to fundamentally rethink the design of manuscript evaluation systems. Rather than replicating workflows optimized for human reviewers, it is crucial to rethink the foundations of effective evaluation and to design AI systems that complement and address the persistent limitations of traditional peer review.
@@
In this work, we propose a shift from replication to redesign. We focus on the decision-making layer of peer review—specifically, mechanisms that can rank submissions and support acceptance decisions. Our approach is intended to complement, not replace, the current peer review process. Specifically, we introduced and explored a novel mechanism that engages LLM agents in pairwise comparisons for more effective manuscript assessment. Rather than assigning absolute scores to individual papers, our approach directly contrasts pairs of submissions to determine which one better meets the multidimensional evaluation criteria. Each LLM agent is tasked with evaluating two manuscripts at a time, and by integrating the results from numerous pairwise comparisons, we can construct a robust overall assessment that more accurately reflects relative quality.
Our results indicate that with proper scaling, this pairwise mechanism demonstrates potential to identify papers of higher academic impact (as measured by future citation count), significantly outperforming paper-level rating approach. However, our analysis also reveals important limitations: Papers preferred by our LLM-based pairwise ranking system tend to exhibit lower topic novelty and greater concentration among a limited number of high-prestige institutions. While these biases highlight critical challenges for the deployment of LLM-based review systems, we view this work as a first step toward exploring fundamentally new paradigms for scalable, equitable, and robust research evaluation.
It is important to note that while we use proxy metrics—such as future citation counts—to evaluate our proposed mechanism, this represents an exploratory and pragmatic step rather than an ideal. The gold standard would be direct human evaluation of how well a mechanism supports the selection and improvement of high-quality research, but such evaluations are costly and difficult to scale [43]. Proxy metrics offer a tractable, though imperfect, way to approximate long-term impact and anticipate how different mechanisms might perform. Ideally, a strong alternative review mechanism should correlate with—but not fully replicate—the outcomes of the current system, offering complementary perspectives and surfacing different strengths. As we show, our proposed pairwise comparison mechanism achieves this balance, aligning with current human judgments while introducing useful differentiation. Crucially, our goal is not to mimic existing outcomes, but to design mechanisms that add value.
Moreover, we believe the strength of human-led peer review lies not only in final decisions, but also in the human interactions it enables—feedback, discussion, and iterative refinement—that help authors strengthen their work. Through these interactions, reviewers and authors also engage in a shared evaluative practice that reinforces norms, builds trust, and helps individuals identify with the scholarly community [63, 19]. Our intention is not to automate or replace this process, but to broaden how we understand and support scholarly evaluation.
@@ 2 Related Work
Peer review, while essential to scientific progress, faces numerous challenges in today’s academic landscape. Reviewers often struggle with overwhelming workloads, further compounded by the absence of meaningful incentives and the inequitable distribution of review requests [7, 31]. In addition, implicit biases can influence evaluation [46, 47, 50], with studies showing that factors like author affiliation or gender may affect acceptance rates [59]. As research output continues to accelerate globally [5], these systemic challenges demand innovative solutions to preserve the integrity and effectiveness of scientific evaluation. Earlier works [48, 57, 56] introduce pairwise comparisons by incorporating author-provided rankings to calibrate noisy reviewer scores. In contrast, we fully replace paper-level scores with large-scale pairwise comparisons performed by LLM agents, without relying on human ratings or author input.
@@
Recent advancements in Large Language Models (LLMs) have sparked considerable interest in their potential to transform academic peer review processes [21, 62, 24, 23]. While existing studies [9, 22, 38] have demonstrated that LLMs can provide valuable feedback for research papers, exhibiting substantial overlap with reviews written by human reviewers, they often struggle with assigning reasonable ratings and making sound decisions, which raises concerns about their reliability for evaluative tasks [26, 58]. For instance, Niu et al. [35] showed that ChatGPT tends to grant acceptances: out of 1558 full paper evaluations, ChatGPT suggested 1243 acceptances compared to only 315 rejections, resulting in an acceptance rate of approximately 79.8%. Similarly, Zhou et al. [60] highlighted that LLMs could generate reasonable aspect scores (e.g. recommendation, soundness, originality) from human reviews but failed when given only the research paper. Lu et al. [27] improved the base LLM’s decision-making process through prompting techniques such as self-reflection [42], few-shot examples [54] and response ensembling [53]. CycleReviewer [55] and DeepReview [61] enhanced LLM-based paper review via fine-tuning on high-quality review dataset. In contrast, Less effort is made to explore new paradigms that fundamentally rethink how LLMs can participate in the academic review process.
@@ 3.2 From Pairwise Comparisons to Rankings
    Pairwise Comparison: Let $\mathcal{S}=\{(i,j):i\neq j\text{ and }i,j\in[N]\}$ be all possible pairs of papers to compare. For each of $M$ agents, we randomly assign one pair $(i,j)\in\mathcal{S}$ to evaluate. Each agent then analyzes both papers and produces a binary preference judgment $y_{ij}\in\{0,1\}$, where $y_{ij}=1$ indicates that the agent prefers paper $i$ over paper $j$, and $y_{ij}=0$ indicates the opposite preference. The full prompt used for pairwise judgment is provided in Supp Figure 7. Denote the collection of all chosen pairs as $\mathcal{A}$, where $|\mathcal{A}|=M$. We collect all comparison results into an observation set $\mathcal{O}=\{(i,j,y_{ij}):(i,j)\in\mathcal{A}\}$, which contains triplets of the compared papers and their corresponding preference judgments from all agents. Since the size of $|\mathcal{S}|$ grows quadratically with $n$, we can only afford to assess a small fraction of all possible pairs, and hence it is challenging to recover the ranking under this scenario. Later in Section 4.2, we empirically examined how ranking quality scales with the number of LLM agents. The results reveal a clear scaling law, and we are able to recover a high-quality ranking with less than 2% samples of all possible pairs.  [^1] Theorem 4 in [34] shows that if $m>12n\log n$ pairs are sampled uniformly at random, then with high probability the estimate $\hat{\theta}$ satisfies $\lVert\hat{\theta}-\theta^{*}\rVert=O\!\left(n\sqrt{\tfrac{\log n}{m}}\right)$.
@@ 4.2 Agent Scaling Boosts the Performance of GPT Ranking System
Figure 2: Scaling of average citation counts for accepted papers by GPT ranking system with increasing pairwise comparisons. We introduce two variants of our GPT ranking system: a basic version without area control, and one with area control, which ensures that the distribution of accepted papers across primary research areas matches that of human peer review. These two variants are compared against three baselines: (1) human acceptance decisions, (2) random acceptance based on the conference acceptance rate, and (3) GPT rating system (acceptance depends on the rating synthesized by a meta-reviewer from three independent GPT reviewers). As the number of comparisons increases, both variants steadily improve and approach human‐level performance. At maximum scale, the GPT ranking system without area control achieves 20.00 average citations, while the version with area control reaches 18.97 average citations. These compare favorably to human acceptance decisions (19.36 average citations), and significantly outperform both random acceptance (8.56 average citations) and the GPT rating system baseline (11.41 average citations).
@@ 4.3 Discriminative Capability of GPT ranking system in Research Evaluation
Figure 3: Comparison of mean citation counts across multiple AI conferences under two‐fold decisions (e.g., accepted vs. rejected, main vs. findings track, oral vs. poster). The results consistently show that higher-tier papers (both GPT-selected and human-selected) receive substantially higher citation counts than lower-tier papers. This implies that our GPT ranking system can effectively distinguish more influential papers, performing comparably to human peer review system in identifying works likely to generate greater impact in the academic community. Error bars represent 95% confidence intervals. *P << 0.05, **P << 0.01, ***P << 0.001, and ****P << 0.0001.
@@ 4.6 Potential Bias against Novel Research Topics
Figure 5: Nearest neighbor distances of papers selected for top-tier acceptance by GPT ranking system versus human reviewers. Higher distances indicate greater novelty. Human-selected papers (blue) consistently show higher nearest neighbor distances than GPT-selected papers (orange) across all conferences. The differences are statistically significant (p < 0.05) for ICLR 2023 Notable-top-5%, ICLR 2024 Oral, and EMNLP 2023 Main Track. For CoRL 2023 Oral and NeurIPS 2023 Oral, we did not observe statistically significant differences, which may be due to smaller sample sizes.
@@ A.1 Limitations
In this study, we rely on citation counts as a proxy for academic impact. Although citation metrics are widely used and provide a quantifiable measure of influence, they are affected by numerous factors beyond quality, such as the visibility of research communities, prevailing research trends, and established reputations [25, 4, 51, 11]. More critically, papers selected by human reviewers inherently receive more visibility through conference presentations and proceedings. The fact that our GPT ranking system identified papers with comparable citation impact despite this disadvantage demonstrates a certain level of effectiveness in our approach. Nevertheless, this circularity problem makes it difficult to establish a truly independent measure of quality. To further verify the robustness of our proposed method, we conducted experiments using alternative measures. Specifically, we measured the Spearman correlation between our system’s output and human review scores. In both ICLR 2023 and ICLR 2024, the BT scores produced by our mechanism show moderate correlation with the average human ratings—27% and 24%, respectively. This indicates that our system aligns with signals in the current review process, while not fully replicating its outcomes. We believe this complementary nature is essential for identifying potential improvements to existing peer review systems.
Furthermore, while we exclusively used GPT-4o mini for our large-scale experiments due to its balance of performance and computational cost, we acknowledge that there are other diverse LLMs capable of academic assessment, and they might yield substantially different results. A more comprehensive evaluation across multiple models would provide better insight into the generalizability of our method and findings. In Appendix C.4, we tested two additional models beyond GPT-4o mini: Gemini 2.0 Flash and Claude-3-Haiku-20240307. We scaled the number of comparisons to over $10^{6}$ and found that, in both cases, our pairwise comparison framework performs significantly better than the GPT rating system baseline. Specifically, papers selected by Gemini 2.0 Flash receive an average of $18.3$ citations (vs. $11.4$), while those chosen by Claude-3-Haiku average $16.8$ (vs. $11.4$). These results suggest that the advantage of our approach is robust across different LLMs.
Finally, while our current work constructs a single global ranking, future extensions could explore personalized or multi-objective evaluation systems that explicitly account for epistemic diversity and evolving community goals.
@@ C.1 Discriminative Capability of GPT ranking system in Research Evaluation
Figure 8: Comparison of mean citation counts across multiple AI conferences under fine-grained decisions in accepted papers. The results consistently show that higher-tier papers (both GPT-selected and human-selected) receive overall higher influence than lower-tier papers. Error bars represent 95% confidence intervals. *P << 0.05, **P << 0.01, ***P << 0.001, and ****P << 0.0001.
@@ C.3 Different Acceptance Patterns across Research Areas
C.4 Using other LLMs as agents
Figure 11: Scaling of average citation counts for accepted papers by Claude ranking system with increasing pairwise comparisons. We observe a similar temporal scaling pattern as in the GPT ranking system, with citation counts increasing as the number of pairwise comparisons grows.
Figure 12: Scaling of average citation counts for accepted papers by Gemini ranking system with increasing pairwise comparisons. We observe a similar temporal scaling pattern as in the GPT ranking system, with citation counts increasing as the number of pairwise comparisons grows.